# AHA! Animating Human Avatars in Diverse Scenes with Gaussian Splatting

### v1

@@ 4 Experiments
For rendering evaluation, we present two modified mesh-based baselines (Baseline A and B) and evaluate with two evaluation protocols (I-human and II-automated). For further evaluation on motion quality and ablations please see supp mat
@@ 4.1 Rendering Evaluation
Baseline A: For Baseline A we collect monocular videos from DL3DV Ling et al. (2024); each scene is reconstructed twice (once as 3DGS, once as a mesh using dense VGGT reconstruction Wang et al. (2025)) so that comparisons are within-scene. Using the meshes obtained using dense VGGT reconstruction, we again use Zhao et al. (2023) to generate SMPL-X parameters. Then we use these parameters naively to pose human Gaussians (Sec. 3.1) in the 3D scene and render the composited scene and human Gaussians using 3DGS. Note we do not perform any refinement. Furthermore note that the scene mesh is only used for motion synthesis but for rendering we use the 3DGS scene reconstruction and the posed human Gaussians. Baseline A is designed to show that a naive baseline that composites human and scenes Gaussians does not work out-of-the-box for monocular videos and hence provides further motivation for our algorithm. For baseline A evaluations, we render synchronized camera trajectories per pair (identical poses, FoV, and exposure).
Baseline B: In this experiment we aim to evaluate the rendering quality of a strong mesh-based baseline vs our 3DGS based algorithm. We use the highest quality existing mesh based 3D scenes from the Replica Straub et al. (2019) dataset. In the Replica Scene we use the framework in Zhao et al. (2023) to synthesize motion. Then we use a rigged scan from RenderPeople (along with its texture map) RenderPeople animated with the synthesized motion parameters in the Replica scene to generate the final renderings. For rendering our videos we use scenes from the Supersplat library and Avatars from Avatarrex Zheng et al. (2023) dataset. This experiment aims to evaluate the highest quality rendering of a mesh based rendering vs a highest quality 3DGS renderings for the specific setting of human-scene animation. Note comparisons are not within scene. Here we also want to acknowledge that for 3DGS animatable Avatars we require multiview video while a mesh can be reconstructed using multiview images. However we believe that for different use-cases, users would be willing to make the tradeoff for higher rendering quality.
@@
Evaluation Protocol II: VLM-based pairwise judgment We use two strong vision–language model (VLM): GPT-5 OpenAI (2025) and Gemini2.5 Comanici et al. (2025) as paired comparators between still renderings. For each video pair, we uniformly sample 10 frames per method, form matched pairs at the same timestamps, and query the VLM with ”which image looks more photoreal?”. The VLM outputs a ternary judgment {Left wins, Right wins, Tie}; we compute a VLM win rate as the percentage of non-tied pairs favoring Ours. We randomize image order and prevent leakage by removing textual overlays and metadata.
@@
Results Baseline A (Same monocular video) On the within-scene comparison, Ours outperforms the mesh baseline in both human and VLM judgments (Tables 23). Note that the scene meshes reconstructed using VGGT often exhibit blocky strucutres, blocked paths and hence are not suitable for motion synthesis, while our algorithm directly operates in Gaussian Space and doesnt suffer from the same problems. For detailed failure modes see Supp Mat.
Baseline B (Replica Scene), HQ vs HQ. Across both comparisons Ours is preferred by humans and by the VLM comparator (Tables 23) - thus clearly underscoring the central premise of the paper - that neural scene representations yield better rendering quality for human-scene interaction compared to existing mesh based representations.
@@ 4.2 Qualitative Results and Free viewpoint rendering of Edited Videos
In Fig. 2, we show results for diverse scenes from the SuperSplat library, Scannet scenes with Avatars from BEHAVE (sparse only 4 cameras), DNA-Rendering Cheng et al. (2023), AvatarX datasets. In Fig. 3, we demonstrate that our method works for monocular RGB scene videos and allows for free viewpoint rendering of videos edited with geometry consistent placement of animated humans in the scene. For more results plz see supp video.
@@ 5 Conclusion
We have presented, to the best of our knowledge, the first method to synthesize human interactions in diverse 3D environments using 3D Gaussian Splatting (3DGS) as the underlying 3D representation. Our results suggest that neural rendering is now mature enough to function as a practical component in end-to-end 3D human–scene animation pipelines, bridging previously disjoint lines of work in human-scene animation and neural rendering. Crucially, our pipeline operates on scenes reconstructed from monocular RGB video and allows for applications such as monocular RGB geometry consistent video editing. We believe this framing and evidence open new research directions at the intersection of human animation, scene understanding, and neural rendering. Despite this progress, our pipeline has several limitations. First, complex and rapidly changing illumination can cause rendering artifacts and imperfect relighting. Second, we do not enforce full physics-based constraints, which can yield interactions that look plausible yet violate contact, stability, or momentum conservation. Third, the range of interaction types is limited; highly dexterous manipulation and long-horizon, multi-contact behaviors remain challenging. Fourth, we assume access to multiview videos of a human performing diverse actions. Addressing these issues suggests several promising directions: stronger lighting estimation and inverse rendering; incorporation of differentiable or learned physics priors for contact and dynamics; expansion to richer, longer, and multi-person interactions; and investigating how to reconstruct avatars from monocular video that generalize to out-of-distribution poses. We hope this work provides a foundation for scalable, video-native human–scene animation pipelines and catalyzes advances in data, models, and evaluation for interactive 3D human animation.

### v2

@@ 4 Experiments
For rendering evaluation, we present two modified mesh-based baselines (Baseline A and B) and evaluate with two evaluation protocols (I-human and II-automated). For further evaluation on motion quality and ablations please see the supplementary.
@@ 4.1 Rendering Evaluation
Baseline A: For Baseline A we collect monocular videos from DL3DV (ling2024dl3dv); each scene is reconstructed twice (once as 3DGS, once as a mesh using dense VGGT reconstruction (wang2025vggt)) so that comparisons are within-scene. Using the meshes obtained using dense VGGT reconstruction, we again use (zhao2023dimos) to generate SMPL-X parameters. Then we use these parameters naively to pose human Gaussians (Sec. 3.1) in the 3D scene and render the composited scene and human Gaussians using 3DGS. Note we do not perform any refinement. Furthermore note that the scene mesh is only used for motion synthesis but for rendering we use the 3DGS scene reconstruction and the posed human Gaussians. Baseline A is designed to show that a naive baseline that composites human and scenes Gaussians does not work out-of-the-box for monocular videos and hence provides further motivation for our algorithm. For baseline A evaluations, we render synchronized camera trajectories per pair (identical poses, FoV, and exposure).
Baseline B: In this experiment we aim to evaluate the rendering quality of a strong mesh-based baseline vs our 3DGS based algorithm. We use the highest quality existing mesh based 3D scenes from the Replica (replica19arxiv) dataset. In the Replica Scene we use the framework in (zhao2023dimos) to synthesize motion. Then we use a rigged scan from RenderPeople (along with its texture map) (RenderPeople) animated with the synthesized motion parameters in the Replica scene to generate the final renderings. For rendering our videos we use scenes from the Supersplat library and Avatars from Avatarrex (zheng2023avatarrex) dataset. This experiment aims to evaluate the highest quality rendering of a mesh based rendering vs a highest quality 3DGS renderings for the specific setting of human-scene animation. Note comparisons are not within scene.
Here we want to highlight that finding the same scene+human combination for both our method and Baseline B would be difficult - because the data capture pipelines for 1) 3DGS vs mesh scene 2) 3DGS vs mesh human are fundamentally different. RenderPeople uses a rig of 250 DSLR cameras to capture a 3D human mesh - while the AvatarX dataset only uses 16 cameras and Behave dataset uses only 4. For 3DGS Avatar reconstruction motion of about 120 seconds inside the multiview camera setup is required while for a mesh capture only one frame is required. The way 3D scenes are reconstructed for Replica (mesh) and SuperSplat (3DGS) are also fundamentally different - the Replica reconstruction uses depth while some Supersplat scenes use lidar. Additionally the Replica original images are not available so we can’t reconstruct a 3DGS splat for Replica scenes of the same quality as the ones on SuperSplat. As such for this particular baselines comparisons are not within scene.
We also want to acknowledge that for 3DGS animatable Avatars we require multiview video while a mesh can be reconstructed using multiview images. However we believe that for different use-cases, users would be willing to make the tradeoff for higher rendering quality.
@@
Evaluation Protocol II: VLM-based pairwise judgment We use two strong vision–language model (VLM): GPT-5 (openai2025gpt5) and Gemini2.5 (comanici2025gemini25pushingfrontier) as paired comparators between still renderings. For each video pair, we uniformly sample 10 frames per method, form matched pairs at the same timestamps, and query the VLM with “which image looks more photoreal?”. The VLM outputs a ternary judgment {Left wins, Right wins, Tie}; we compute a VLM win rate as the percentage of non-tied pairs favoring Ours. We randomize image order and prevent leakage by removing textual overlays and metadata.
For the pairwise VLM study, the VLM is instructed to ignore artistic style and focus on physical plausibility: “which image looks more photoreal? Consider geometry (straight lines, depth cues), materials (BRDF, speculars), lighting/shadows, and absence of artifacts (flicker, halos, floaters).
@@
Results Baseline A (Same monocular video) On the within-scene comparison (Fig. 5), Ours outperforms the mesh baseline in both human and VLM judgments (Tables 23). Note that the scene meshes reconstructed using VGGT often exhibit blocky strucutres, blocked paths and hence are not suitable for motion synthesis, while our algorithm directly operates in Gaussian Space and doesnt suffer from the same problems. For detailed failure modes see Supp Mat.
Baseline B (Replica Scene), HQ vs HQ. Across both evaluations (Fig. 5) Ours is preferred by humans and by the VLM comparator (Tables 23) - thus clearly underscoring the central premise of the paper - that neural scene representations yield better rendering quality for human-scene interaction compared to existing mesh based representations.
@@ 4.2 Qualitative Results and Free viewpoint rendering of Edited Videos
In Fig. 3, we show results for diverse scenes from the SuperSplat library, Scannet scenes with Avatars from BEHAVE (sparse only 4 cameras), DNA-Rendering (cheng2023dnarendering), Avatarrex (zheng2023avatarrex) datasets. In Fig. 4, we demonstrate that our method works for monocular RGB scene videos and allows for free viewpoint rendering of videos edited with geometry consistent placement of animated humans in the scene. For more results please see the supplementary.
@@ 5 Conclusion
We have presented, to the best of our knowledge, the first method to synthesize human interactions in diverse 3D environments using 3D Gaussian Splatting (3DGS) as the underlying 3D representation. Our results suggest that neural rendering is now mature enough to function as a practical component in end-to-end 3D human–scene animation pipelines, bridging previously disjoint lines of work in human-scene animation and neural rendering. Crucially, our pipeline operates on scenes reconstructed from monocular RGB video and allows for applications such as monocular RGB geometry consistent video editing. We believe this framing and evidence open new research directions at the intersection of human animation, scene understanding, and neural rendering.
Limitations. Despite this progress, our pipeline has several limitations. First, complex and rapidly changing illumination can cause rendering artifacts and imperfect relighting. Second, we do not enforce full physics-based constraints, which can yield interactions that look plausible yet violate contact, stability, or momentum conservation. Third, the range of interaction types is limited; highly dexterous manipulation and long-horizon, multi-contact behaviors remain challenging. Fourth, we assume access to multiview videos of a human performing diverse actions.
Outlook. Addressing these issues suggests several promising research directions: integrating stronger lighting estimation and inverse rendering, incorporating differentiable or learned physics priors for contact and dynamics, expanding the interaction vocabulary to richer, longer, and multi-person scenarios and investigation how Avatars that generalize to Out-of-distribution poses can be reconstructed from monocular videos. We hope this work provides a foundation for scalable, video-native human–scene animation pipelines and catalyzes further advances in data, models, and evaluation for interactive 3D human animation.